# CogFlow: Bridging Perception and Reasoning through Knowledge Internalization for Visual Mathematical Problem Solving

### v1

@@ 1 Introduction
Multimodal large language models (MLLMs) are rapidly advancing and have been applied across various vision–language applications (peng2024grounding; xu2025mc; yue2024mmmu). However, existing MLLMs continue to struggle with challenging visual mathematical problems, resulting in low answer accuracy and inconsistent reasoning chains. Some early attempts (VL-Rethinker; VLM-R1) adopt a one-step reasoning framework that directly interleaves visual perception with reasoning in an unstructured manner, often resulting in both perceptual and reasoning errors. Another line of work (MathFlow; DVLR) follows a decoupled reasoning pipeline that explicitly separates the perception and reasoning parts, with the former focusing on visual recognition and the latter responsible for subsequent inference. Yet in practice, we observe that such a pipeline often suffers from the reasoning drift issue, i.e., it tends to yield illogical or unwarranted reasoning steps that disregard perceptual evidence (see Figure 1). These observations motivate the development of a new approach that not only achieves robust fine-grained recognition of mathematical visual elements (e.g., diagrammatic primitives and symbols) but also faithfully incorporates extracted visual cues into subsequent reasoning.
@@
Figure 1: The one-step reasoning framework (e.g., VLM-R1 (VLM-R1)) often yields suboptimal results, while the decoupled reasoning pipeline (e.g., MathFlow (MathFlow)) enhances perception yet still yields illogical reasoning steps that disregard visual evidence. In contrast, CogFlow adopts a cognitive-inspired three-stage framework to effectively mitigate reasoning drift.
@@
Unlike prior approaches (vcar; MathFlow; DVLR) that decouple perception from reasoning trajectories (rlhf) and enhance it with tailored tasks, CogFlow first integrates perception enhancement into a unified reinforcement learning (RL) framework through Synergistic Visual Rewards (SynVRs), enabling dynamic perception–reasoning interaction and improving generalization. Specifically, SynVRs complementarily optimize the model from two distinct perspectives: (1) a Visual Parameterized Reward (VPR) encoding normalized primitives (i.e., points, lines, and circles) and calculating the Euclidean distance in a parameter space for precise and interpretable measurement; (2) a Visual Semantic Reward (VSR) that extracts semantic embeddings (fg-clip) from re-rendered images (derived from textual perception outputs) and measures the cosine distance in a semantic space to capture holistic style and layout consistency. Together, SynVRs ensure both local geometric fidelity and global perceptual coherence, forming trustworthy visual cues that serve as the foundation for effective visual mathematical reasoning.
@@
Notably, despite progress in perception enhancement, all prior efforts (vcar; DVLR; ovr) remain confined to accurate extraction of mathematical information from diagrams, while ignoring a key question: are the extracted visual cues properly and faithfully integrated into subsequent reasoning? As illustrated in Figure 1(b), our empirical findings reveal a typical reasoning drift issue (i.e., the reasoning stage in existing methods often deviates from perceptual results), leading to reasoning chains that appear coherent yet conflict with the underlying visual evidence. To prevent such drift and improve interpretability, CogFlow utilizes a Knowledge Internalization Reward (IntlzR) that bridges the perception and reasoning stages by encouraging the model to generate structured and reasoning-ready outputs (i.e., knowledge-internalized representations (cognition4)) as a more reliable foundation for subsequent reasoning. Specifically, we curate positive trajectories integrating perception and reasoning processes with explicit internalization of perception primitives, and further derive five typical negative trajectories. Training with these trajectories enables the reward model to evaluate each response according to its fidelity to the internalized representation. IntlzR effectively improves the knowledge internalization stage, thereby reducing hallucinations and improving interpretability and robustness.
@@
In accordance with the hierarchical flow of human reasoning, we further improve multi-step visual reasoning beyond enhanced perception and knowledge internalization. Existing approaches either follow a text-centric RL paradigm (deepseek-r1) that is free from perceptual objectives (chen2025sft; wang2025think), or overlook the structured dependency between perception and reasoning (VLM-R1; VL-Rethinker). To ensure more stable reasoning in the presence of perceptual errors, CogFlow introduces Visual-Gated Policy Optimization (VGPO) strategy that explicitly anchors the reasoning process in perceptual accuracy. In VGPO, a visual gate is designed to adaptively filter perceptual trajectories through perceptual quality assessment, retaining only high-quality ones before subsequent reasoning trajectory generation. If a low-quality perceptual trajectory is filtered out, the model regenerates alternative trajectories to obtain a higher-quality response. Along with the proposed visual gate, VGPO integrates an outcome-supervised Inference Reward (grpo) for optimization, further strengthening multi-step visual reasoning.
@@
To facilitate research, we curate a new MathCog dataset for model training, which contains three subsets and over 120K samples with high-quality perception-reasoning aligned annotations. We conduct extensive experiments on commonly used visual math problem-solving benchmarks (MathFlow; zhang2024mathverse; lu2024mathvista; wemath; logicvista; dynamath) to comprehensively evaluate CogFlow. The results show that CogFlow consistently outperforms state-of-the-art MLLMs with comparable model sizes. Notably, it achieves on-par or even better results compared to advanced closed-source MLLMs with much larger model sizes. The main contributions of this paper can be summarized as follows:
@@ 2 Related Work
Visual Mathematical Reasoning with MLLMs. Solving visual mathematical problems (e.g., geometry diagrams, algebraic plots, and etc) requires both strong reasoning ability and accurate interpretation of visual primitives and symbolic content (yan2024survey; wemath2; SVE-Math-Qwen2.5-7B). Most previous works are dedicated to improving the reasoning process, including chain-of-thought strategies (xu2025llavacotletvisionlanguage; r-cot), tool-aided reasoning (alphageometry; pot), test time scaling (visuothink; vstar), and reinforcement learning (Skyworkr1v2; vlm-r3). Several recent works (DVLR; vcar; wei2025slowperceptionletsperceive) suggest that one of the major bottlenecks in visual mathematical reasoning is inaccurate visual comprehension. They typically decouple perception from reasoning, and strengthen perception either by designing specialized visual encoders (zhang2024mavismathematicalvisualinstruction) or by introducing auxiliary visual tasks (MathFlow). However, prior works ignore a key issue of whether correctly extracted visual cues are indeed faithfully incorporated into subsequent reasoning.
Reinforcement Learning for Multimodal Reasoning. Traditional actor–critic methods, such as proximal policy optimization (ppo1), are computationally expensive. A lightweight alternative is group relative policy optimization (GRPO) (deepseek-r1), which stabilizes advantage estimation using group baselines. While its variants (gspo; DAPO) have been widely explored, GRPO has also been extended to multimodal reasoning (Skyworkr1v2; VLM-R1; RLHF-V). Some extensions introduce hybrid reward formulations augmented with preference signals during training (Skyworkr1v2; medvlm), whereas others propose two-stage multimodal reinforcement learning paradigms, e.g., OVR (ovr). However, existing methods typically lack explicit mechanisms to strengthen the alignment between perception and reasoning, often leading to reasoning that is not firmly grounded in visual content.
@@ 3.1 Forming Trustworthy Perception with Synergistic Visual Rewards
We then compute the cost matrix $\mathcal{C}$ between GT Primitives $\mathcal{G}$ and predicted Primitives $\mathcal{P}$ and apply the Hungarian matching algorithm (hungarian) to obtain the optimal one-to-one matching $\mathcal{H}$ that minimizes the total cost $\mathcal{S}_{\text{VPR}}$. The VPR offers interpretable, geometry-aware semantic supervision that avoids the pitfalls of pixel-level noise or black-box embedding similarity. More details are provided in §C.2.
Capturing Holistic Layout and Style Consistency in Semantic Space. We render the predicted response into an image $\tilde{I}$ and compare it against the ground-truth rendering $I$, using a frozen FG-CLIP encoder (fg-clip) $\phi(\cdot)$ with higher sensitivity to spatial features. The VSR score $\mathcal{S}_{\mathrm{VSR}}$ is the normalized cosine similarity. The higher values indicate closer agreement in global layout and style while preserving fine-grained geometric fidelity.
@@ 3.2 Alleviating Reasoning Drift with Knowledge Internalization Reward
Rewarding Faithful Knowledge Internalization. To address these systematic weaknesses, we introduce the Knowledge Internalization Reward (IntlzR), a trained reward model that evaluates whether each reasoning chain remains faithful to the internalized visual representation, thereby enforcing raw perception as the primary substrate of inference (e.g., reasoning drift). Within CogFlow, we train the IntlzR to enhance internalization explicitly, supported by the curated MathCog-IntlzR (see §C.3 for more details). This dataset is constructed from positive–negative pairs, with each positive matched to five corresponding negatives; positives are sampled from MathCog-SFT, and negatives are synthesized by injecting the five error types described above, providing fine-grained supervision for distinguishing grounded from unanchored reasoning. To better leverage diverse negative signals in IntlzR and adaptively emphasize the most challenging trajectories while improving gradient efficiency and training stability, we adopt Softmax-DPO (softmaxdpo) for optimization:
@@ 4.1 Experimental Setup
Evaluation Benchmarks and Metrics. We conduct experiments on widely used benchmarks (i.e., FlowVerse (MathFlow), MathVerse (zhang2024mathverse), MathVista (lu2024mathvista), WeMath (wemath), LogicVista (logicvista), and DynaMath (dynamath)) with various visual mathematical reasoning tasks and different visual and textual complexity. Following common practice, we measure the accuracy (Acc) of the final answer and assess the reasoning ability from the intermediate reasoning process using the chain-of-thought evaluation (CoT-E).
Baselines. We compare our method with a wide range of closed-source MLLMs, such as GPT series (gpt4o; gpt4v; gpt5), Gemini series (gemini), DouBao-1.5-pro (doubao), GLM-4.5V (glm), Claude (claude-3-5) and Qwen-VL-Plus (bai2023qwenvlversatilevisionlanguagemodel), as well as open-source MLLMs, including InfiMM-Math (infimm), InternVL2.5 (internvl2.5), InternLM-XC2 (internlm), Math-LLaVA (Math-LLaVA), MultiMath-7B (MultiMath), SVE-Math (SVE-Math-Qwen2.5-7B), VLM-R1 (VLM-R1), MathFlow (MathFlow), DVLR (DVLR), VCAR (vcar), VL-Rethinker (VL-Rethinker), ThinkLite-VL (thinklite) and SPHINX-MoE (sphinx). Since some concurrent works have yet to release their models, we report the results provided in their original papers, some of which do not include category-wise performance.
Implementation Details. We initialize CogFlow with Qwen2.5-VL-7B (Qwen2.5-VL) and train it on our curated MathCog dataset. For cold start, we first train CogFlow on MathCog-SFT subset for 2 epochs with a learning rate of $1\times 10^{-5}$ and a batch size of 64. Subsequently, CogFlow is optimized using VGPO on MathCog-RL subset for 1 epoch with a learning rate of $1\times 10^{-6}$ and a batch size of 16. The IntlzR reward model is based on Qwen2.5-VL-3B and trained on MathCog-IntlzR subset for 3 epochs with a learning rate of $7\times 10^{-6}$ and a batch size 64. All our models are trained on 16 NVIDIA A100 GPUs. Please refer to §C for more detailed configures.
@@ 4.3 Ablation Studies
Error Type Analysis. As shown in Figure 9, we employ GPT-5 (gpt5) to classify each response on the FlowVerse benchmark into one of four categories: Perception Error, Knowledge Internalization Error, Reasoning Error, and Correct. The first row covers specialized visual-math MLLMs (MultiMath-7B (MultiMath), SVE-Math-7B (SVE-Math-Qwen2.5-7B)), an GRPO-style model (VLM-R1 (VLM-R1)), the decoupled method (MathFlow-7B (MathFlow)), and GPT-4o (gpt4o); the second row traces our variants from the SFT+GRPO baseline through +SynVRs, +IntlzR, and +VGPO to CogFlow. We observe: (1) GPT-4o’s strong overall performance is driven by superior reasoning, yet its diagram perception lags behind specialized 7B models, indicating that strengthening perception and internalization remains essential; (2) Although methods such as MathFlow-7B and VLM-R1 reduce Perception Error and Reasoning Error to some extent, these methods do not meaningfully alleviate Knowledge Internalization Error; (3) +SynVRs chiefly reduces Perception Errors ($2\%$), confirming that geometry-aware and semantic visual rewards improve perceptual fidelity; (4) +IntlzR primarily mitigates reasoning drift, reducing Knowledge-Internalization Error by $2\%$ relative to the baseline; and (5) when all components are combined, CogFlow minimizes every error type while maximizing the proportion of correct predictions. Furthermore, we conduct a systematic case study (see §D.13) to analyze the error compensation mechanism.

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) are rapidly advancing and have been applied across various vision–language applications (Peng et al., 2024b; Xu et al., 2025b; Yue et al., 2024). However, existing MLLMs continue to struggle with challenging visual mathematical problems, resulting in low answer accuracy and inconsistent reasoning chains. Some early attempts (Wang et al., 2025a; Shen et al., 2025) adopt a one-step reasoning framework that directly interleaves visual perception with reasoning in an unstructured manner, often resulting in both perceptual and reasoning errors. Another line of work (Chen et al., 2025a; Guo et al., 2025c) follows a decoupled reasoning pipeline that explicitly separates the perception and reasoning parts, with the former focusing on visual recognition and the latter responsible for subsequent inference. Yet in practice, we observe that such a pipeline often suffers from the reasoning drift issue, i.e., it tends to yield illogical or unwarranted reasoning steps that disregard perceptual evidence (see Figure 1). These observations motivate the development of a new approach that not only achieves robust fine-grained recognition of mathematical visual elements (e.g., diagrammatic primitives and symbols) but also faithfully incorporates extracted visual cues into subsequent reasoning.
@@
Figure 1: The one-step reasoning framework (e.g., VLM-R1 (Shen et al., 2025)) often yields suboptimal results, while the decoupled reasoning pipeline (e.g., MathFlow (Chen et al., 2025a)) enhances perception yet still yields illogical reasoning steps that disregard visual evidence. In contrast, CogFlow adopts a cognitive-inspired three-stage framework to effectively mitigate reasoning drift.
@@
Unlike prior approaches (Jia et al., 2024; Chen et al., 2025a; Guo et al., 2025c) that decouple perception from reasoning trajectories (Ouyang et al., 2022) and enhance it with tailored tasks, CogFlow first integrates perception enhancement into a unified reinforcement learning (RL) framework through Synergistic Visual Rewards (SynVRs), enabling dynamic perception–reasoning interaction and improving generalization. Specifically, SynVRs complementarily optimize the model from two distinct perspectives: (1) a Visual Parameterized Reward (VPR) encoding normalized primitives (i.e., points, lines, and circles) and calculating the Euclidean distance in a parameter space for precise and interpretable measurement; (2) a Visual Semantic Reward (VSR) that extracts semantic embeddings (Xie et al., 2025) from re-rendered images (derived from textual perception outputs) and measures the cosine distance in a semantic space to capture holistic style and layout consistency. Together, SynVRs ensure both local geometric fidelity and global perceptual coherence, forming trustworthy visual cues that serve as the foundation for effective visual mathematical reasoning.
@@
Notably, despite progress in perception enhancement, all prior efforts (Jia et al., 2024; Guo et al., 2025c; Wei et al., 2025) remain confined to accurate extraction of mathematical information from diagrams, while ignoring a key question: are the extracted visual cues properly and faithfully integrated into subsequent reasoning? As illustrated in Figure 1(b), our empirical findings reveal a typical reasoning drift issue (i.e., the reasoning stage in existing methods often deviates from perceptual results), leading to reasoning chains that appear coherent yet conflict with the underlying visual evidence. To prevent such drift and improve interpretability, CogFlow utilizes a Knowledge Internalization Reward (IntlzR) that bridges the perception and reasoning stages by encouraging the model to generate structured and reasoning-ready outputs (i.e., knowledge-internalized representations (Ryan and Connell, 1989)) as a more reliable foundation for subsequent reasoning. Specifically, we curate positive trajectories integrating perception and reasoning processes with explicit internalization of perception primitives, and further derive five typical negative trajectories. Training with these trajectories enables the reward model to evaluate each response according to its fidelity to the internalized representation. IntlzR effectively improves the knowledge internalization stage, thereby reducing hallucinations and improving interpretability and robustness.
@@
In accordance with the hierarchical flow of human reasoning, we further improve multi-step visual reasoning beyond enhanced perception and knowledge internalization. Existing approaches either follow a text-centric RL paradigm (Guo et al., 2025a) that is free from perceptual objectives (Chen et al., 2025b; Wang et al., 2025b), or overlook the structured dependency between perception and reasoning (Shen et al., 2025; Wang et al., 2025a). To ensure more stable reasoning in the presence of perceptual errors, CogFlow introduces Visual-Gated Policy Optimization (VGPO) strategy that explicitly anchors the reasoning process in perceptual accuracy. In VGPO, a visual gate is designed to adaptively filter perceptual trajectories through perceptual quality assessment, retaining only high-quality ones before subsequent reasoning trajectory generation. If a low-quality perceptual trajectory is filtered out, the model regenerates alternative trajectories to obtain a higher-quality response. Along with the proposed visual gate, VGPO integrates an outcome-supervised Inference Reward (Shao et al., 2024) for optimization, further strengthening multi-step visual reasoning.
@@
To facilitate research, we curate a new MathCog dataset for model training, which contains three subsets and over 120K samples with high-quality perception-reasoning aligned annotations. We conduct extensive experiments on commonly used visual math problem-solving benchmarks (Chen et al., 2025a; Zhang et al., 2024; Lu et al., 2024; Qiao et al., 2025a; Xiao et al., 2024; Zou et al., 2025) to comprehensively evaluate CogFlow. The results show that CogFlow consistently outperforms state-of-the-art MLLMs with comparable model sizes. Notably, it achieves on-par or even better results compared to advanced closed-source MLLMs with much larger model sizes. The main contributions of this paper can be summarized as follows:
@@ 2 Related Work
Visual Mathematical Reasoning with MLLMs. Solving visual mathematical problems (e.g., geometry diagrams, algebraic plots, and etc) requires both strong reasoning ability and accurate interpretation of visual primitives and symbolic content (Yan et al., 2025; Qiao et al., 2025b; Zhang et al., 2025b). Most previous works are dedicated to improving the reasoning process, including chain-of-thought strategies (Xu et al., 2025a; Deng et al., 2024), tool-aided reasoning (Trinh et al., 2024; Chen et al., 2023b), test time scaling (Wang et al., 2025e; Hosseini et al., 2024), and reinforcement learning (Wang et al., 2025c; Jiang et al., 2025). Several recent works (Guo et al., 2025c; Jia et al., 2024; Wei et al., 2024) suggest that one of the major bottlenecks in visual mathematical reasoning is inaccurate visual comprehension. They typically decouple perception from reasoning, and strengthen perception either by designing specialized visual encoders (Zhang et al., 2025a) or by introducing auxiliary visual tasks (Chen et al., 2025a). However, prior works ignore a key issue of whether correctly extracted visual cues are indeed faithfully incorporated into subsequent reasoning.
Reinforcement Learning for Multimodal Reasoning. Traditional actor–critic methods, such as proximal policy optimization (Yu et al., 2022), are computationally expensive. A lightweight alternative is group relative policy optimization (GRPO) (Guo et al., 2025a), which stabilizes advantage estimation using group baselines. While its variants (Zheng et al., 2025; Yu et al., 2025) have been widely explored, GRPO has also been extended to multimodal reasoning (Wang et al., 2025c; Shen et al., 2025; Zhou et al., 2025). Some extensions introduce hybrid reward formulations augmented with preference signals during training (Wang et al., 2025c; Pan et al., 2025), whereas others propose two-stage multimodal reinforcement learning paradigms, e.g., OVR (Wei et al., 2025). However, existing methods typically lack explicit mechanisms to strengthen the alignment between perception and reasoning, often leading to reasoning that is not firmly grounded in visual content.
@@ 3.1 Forming Trustworthy Perception with Synergistic Visual Rewards
We then compute the cost matrix $\mathcal{C}$ between GT Primitives $\mathcal{G}$ and predicted Primitives $\mathcal{P}$ and apply the Hungarian matching algorithm (Kuhn, 1955) to obtain the optimal one-to-one matching $\mathcal{H}$ that minimizes the total cost $\mathcal{S}_{\text{VPR}}$. The VPR offers interpretable, geometry-aware semantic supervision that avoids the pitfalls of pixel-level noise or black-box embedding similarity. More details are provided in §C.2.
Capturing Holistic Layout and Style Consistency in Semantic Space. We render the predicted response into an image $\tilde{I}$ and compare it against the ground-truth rendering $I$, using a frozen FG-CLIP encoder (Xie et al., 2025) $\phi(\cdot)$ with higher sensitivity to spatial features. The VSR score $\mathcal{S}_{\mathrm{VSR}}$ is the normalized cosine similarity. The higher values indicate closer agreement in global layout and style while preserving fine-grained geometric fidelity.
@@ 3.2 Alleviating Reasoning Drift with Knowledge Internalization Reward
Rewarding Faithful Knowledge Internalization. To address these systematic weaknesses, we introduce the Knowledge Internalization Reward (IntlzR), a trained reward model that evaluates whether each reasoning chain remains faithful to the internalized visual representation, thereby enforcing raw perception as the primary substrate of inference (e.g., reasoning drift). Within CogFlow, we train the IntlzR to enhance internalization explicitly, supported by the curated MathCog-IntlzR (see §C.3 for more details). This dataset is constructed from positive–negative pairs, with each positive matched to five corresponding negatives; positives are sampled from MathCog-SFT, and negatives are synthesized by injecting the five error types described above, providing fine-grained supervision for distinguishing grounded from unanchored reasoning. To better leverage diverse negative signals in IntlzR and adaptively emphasize the most challenging trajectories while improving gradient efficiency and training stability, we adopt Softmax-DPO (Chen et al., 2024a) for optimization:
@@ 4.1 Experimental Setup
Evaluation Benchmarks and Metrics. We conduct experiments on widely used benchmarks (i.e., FlowVerse (Chen et al., 2025a), MathVerse (Zhang et al., 2024), MathVista (Lu et al., 2024), WeMath (Qiao et al., 2025a), LogicVista (Xiao et al., 2024), and DynaMath (Zou et al., 2025)) with various visual mathematical reasoning tasks and different visual and textual complexity. Following common practice, we measure the accuracy (Acc) of the final answer and assess the reasoning ability from the intermediate reasoning process using the chain-of-thought evaluation (CoT-E).
Baselines. We compare our method with a wide range of closed-source MLLMs, such as GPT series (OpenAI, 2024; 2023; 2025), Gemini series (Team et al., 2025), DouBao-1.5-pro (Guo et al., 2025b), GLM-4.5V (Hong et al., 2025), Claude (Anthropic, 2024) and Qwen-VL-Plus (Bai et al., 2023), as well as open-source MLLMs, including InfiMM-Math (Han et al., 2025), InternVL2.5 (Chen et al., 2024b), InternLM-XC2 (Dong et al., 2024), Math-LLaVA (Shi et al., 2024), MultiMath-7B (Peng et al., 2024a), SVE-Math (Zhang et al., 2025b), VLM-R1 (Shen et al., 2025), MathFlow (Chen et al., 2025a), DVLR (Guo et al., 2025c), VCAR (Jia et al., 2024), VL-Rethinker (Wang et al., 2025a), ThinkLite-VL (Wang et al., 2025d) and SPHINX-MoE (Liu et al., 2024a). Since some concurrent works have yet to release their models, we report the results provided in their original papers, some of which do not include category-wise performance.
Implementation Details. We initialize CogFlow with Qwen2.5-VL-7B (Bai et al., 2025) and train it on our curated MathCog dataset. For cold start, we first train CogFlow on MathCog-SFT subset for 2 epochs with a learning rate of $1\times 10^{-5}$ and a batch size of 64. Subsequently, CogFlow is optimized using VGPO on MathCog-RL subset for 1 epoch with a learning rate of $1\times 10^{-6}$ and a batch size of 16. The IntlzR reward model is based on Qwen2.5-VL-3B and trained on MathCog-IntlzR subset for 3 epochs with a learning rate of $7\times 10^{-6}$ and a batch size 64. All our models are trained on 16 NVIDIA A100 GPUs. Please refer to §C for more detailed configures.
@@ 4.3 Ablation Studies
Error Type Analysis. As shown in Figure 9, we employ GPT-5 (OpenAI, 2025) to classify each response on the FlowVerse benchmark into one of four categories: Perception Error, Knowledge Internalization Error, Reasoning Error, and Correct. The first row covers specialized visual-math MLLMs (MultiMath-7B (Peng et al., 2024a), SVE-Math-7B (Zhang et al., 2025b)), an GRPO-style model (VLM-R1 (Shen et al., 2025)), the decoupled method (MathFlow-7B (Chen et al., 2025a)), and GPT-4o (OpenAI, 2024); the second row traces our variants from the SFT+GRPO baseline through +SynVRs, +IntlzR, and +VGPO to CogFlow. We observe: (1) GPT-4o’s strong overall performance is driven by superior reasoning, yet its diagram perception lags behind specialized 7B models, indicating that strengthening perception and internalization remains essential; (2) Although methods such as MathFlow-7B and VLM-R1 reduce Perception Error and Reasoning Error to some extent, these methods do not meaningfully alleviate Knowledge Internalization Error; (3) +SynVRs chiefly reduces Perception Errors ($2\%$), confirming that geometry-aware and semantic visual rewards improve perceptual fidelity; (4) +IntlzR primarily mitigates reasoning drift, reducing Knowledge-Internalization Error by $2\%$ relative to the baseline; and (5) when all components are combined, CogFlow minimizes every error type while maximizing the proportion of correct predictions. Furthermore, we conduct a systematic case study (see §D.13) to analyze the error compensation mechanism.